OpenStudy (anonymous):

2x+4y=5 find the slope. please explain.

OpenStudy (anonymous):

for the slope we have to reduce it to form of y=mx+c hence here 4y=-2x +5 or y=-2x/4 +5/4 or y=(-1/2) x +1.25 hence slope is -1/2

OpenStudy (anonymous):

how about 4x-3y+3